A passenger train travels 295 mi in the same amount of time it takes a freight train to travel 225 mi. The rate of the passenger train is 14mph greater than the rate of the freight train. Find the rate of each train.
Answer by mananth(16949)

speed of freight train =x
passanger train speed = x+14
...
passanger train distance = 295
freight train distance = 225
...
Time = distance/speed
both take same time
295/x+14 = 225/x
295x=225(x+14)
295x-225x= 3150
70x=3150
x= 45 mph speed of freight train
passanger train speed = 45+14 = 59 mph
...
m.ananth@hotmail.ca